enthalpy of the steam at the entry of condenser is 200 kilo joule enthalpy of water at the exit of condenseris 100 kj what will be the work done by condensor


1
Expert's answer
2021-07-19T15:37:31-0400

Enthalpy

H=U+PV

dH=dU+PdV

"H_2-H_1=dU+PdV"

Put value

"200-100=dU+PdV"

Isothermal process

dU=0

Put dU value


"PdV=H_2-H_1=200-100=100KJ"

"dW=PdV"

Then

dW=PdV=100KJ

Work done by condensor

dW=100kJ